OpenStudy (anonymous):

Let logP/N=3 and logM/N=6 what is the relationship between P and M? Help??

OpenStudy (precal):

is the P/N in ( )? Log (P/N)=3

OpenStudy (anonymous):

No clue, i typed it just how it was asked.

OpenStudy (precal):

well sorry in that case I have no idea, if it did then I could use log properties.....

OpenStudy (anonymous):

Thanks so much anyway! I appreciate it.

OpenStudy (ash2326):

Subtract the two \[\log \frac{M}{N}-\log \frac{M}{N}=6-3\] Can you simplify this?

OpenStudy (ash2326):

oops \[\log \frac{M}{N}-\log \frac{P}{N}=6-3\]

OpenStudy (anonymous):

I'm completely lost on this problem.

OpenStudy (ash2326):

\[\log a-\log b=\log \frac ab\] Try to use this here